OpenStudy (anonymous):

-1/3x + y = –1 y = 4 + 1/3x

OpenStudy (anonymous):

are you supposed to find where they intersect?

OpenStudy (anonymous):

the first one can be rewritten as \(y=\frac{1}{3}x-1\) so you can set \[\frac{1}{3}x-1=4+\frac{1}{3}x\] and try to solve for \(x\) you can see that the slopes are both \(\frac{1}{3}\) so the lines to not intersect